OpenStudy (anonymous):

The probability of drawing a yellow marble from a bag is 15%. What is the probability of pulling a marble that is not yellow from the bag?

OpenStudy (anonymous):

85%?

OpenStudy (♪chibiterasu):

100 - 15 = 85

OpenStudy (triciaal):

percent means per 100 yellow = 15/100 so not yellow = 100/100-15/100 = 85/100 = 85% as above